OpenStudy (millsemily):

The value of y varies directly with x. What is k the constant of variation for this function, and what is y for the given value of x? y= -1/2 when x=6. Find y when x=-3 Can someone explain this as they solve? Because I have a lot more of them, and I'd like to do the rest on my own :) Thanks!

rishavraj (rishavraj):

as it varies directly so y = kx

rishavraj (rishavraj):

get "k" from above equation by plugging given value of y and x

imqwerty (imqwerty):

so y varies directly with x y ∝ x now removing the proportionality sign '∝' and putting constant k we get y=kx we know when y=-1/2 , x =6 so -1/2 =k(6) so k=-1/12

OpenStudy (millsemily):

And the other is 1/4 , right?

imqwerty (imqwerty):

yea

OpenStudy (millsemily):

Thanks, wasn't as hard as I was making it xD

imqwerty (imqwerty):

( ͡° ͜ʖ ͡°)